OpenStudy (anonymous):

Help please! can someone show me how to get from [(3n - 1) / (5n+2)] to [((3/5)(5n+2) - (6/5) - 1) / (5n+2)] ?

OpenStudy (mayankdevnani):

is this your question ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

help me please T.T

OpenStudy (anonymous):

its says to use the 'juggling' method, but i dont know what is juggling method

OpenStudy (mayankdevnani):

first solve this :- \[\large \bf \frac{3}{5} \times (5n+2)=\frac{15n+6}{5}\]

OpenStudy (anonymous):

3n + (6/5)

OpenStudy (mayankdevnani):

then, \[\large \bf \frac{15n+6}{5}-\frac{6}{5}-1\] Take LCM 5, \[\large \bf \frac{15n+6-6-5}{5}\] we get, \[\large \bf \frac{15n-5}{5}=\frac{15n}{5}-\frac{5}{5}=3n-1\] After all,we get \[\large \bf \frac{3n-1}{5n+2}=\color{red}{Answer}\]

OpenStudy (mayankdevnani):

understood? @HatcrewS

OpenStudy (anonymous):

can i know what do you use to type in fraction form ?

OpenStudy (anonymous):

uhm.. still dont quite get it.. how does the (3/5) come about in the first palce

OpenStudy (anonymous):

place*

OpenStudy (mayankdevnani):

yeah,why not? use this :- `frac{x}{y}`=\(\large \bf \frac{x}{y}\) { Remember:- Type this in equation editor }

OpenStudy (anonymous):

uhm.. sir how do you get [(3n - 1) / (5n+2)] to [((3/5)(5n+2) - (6/5) - 1) / (5n+2)] ? you are solving it the other way round @@

OpenStudy (mayankdevnani):

in which step you stucked ? @HatcrewS

OpenStudy (anonymous):

i dont even know how to begin

OpenStudy (mayankdevnani):

first of all, do you get this ?

OpenStudy (anonymous):

\[\frac{3n-1}{5n-2} \] to \[\frac{\frac{3}{5}(5n+2) - \frac{6}{5} - 1}{5n+2}\] to \[\frac{3}{5} - \frac{11}{5(5n+2)}\]

OpenStudy (anonymous):

sorry its \[\frac{3n-1}{5n+2}\]*

OpenStudy (mayankdevnani):

your last step is wrong..first solve numerator first and then go to denominator

OpenStudy (mayankdevnani):

@HatcrewS just concentrate on my work...

OpenStudy (mayankdevnani):

please read my last 3 rd comment....do you get this or not ?

OpenStudy (anonymous):

i understand the working, but i dont know how do you get to have that equation

OpenStudy (anonymous):

where does the \[\frac{3}{5}\] comes about ?

OpenStudy (mayankdevnani):

i only took this equation under red line

OpenStudy (anonymous):

the steps that i've written are the steps given in the textbook, i just dont get it how it works

OpenStudy (mayankdevnani):

i solved the bracket first

OpenStudy (mayankdevnani):

@HatcrewS understood?

OpenStudy (anonymous):

no. i need to understand your working. why do you bring (5n+2) to the numerator and where does the (3/5) comes from ???

OpenStudy (mayankdevnani):

go through your question, just find out 3/5

OpenStudy (anonymous):

at first you only see this \[\frac{3n-1}{5n+2}\] what are the steps or mathematical method do you use to make it become like this \[\frac{\frac{3}{5}(5n+2) - \frac{6}{5} - 1}{5n+2} \]

OpenStudy (mayankdevnani):

no..you are saying opposite ....

OpenStudy (mayankdevnani):

its so hard to get this 2 nd equation from 1 st equation

OpenStudy (anonymous):

exactly... thats what ive been talking about... its from 1st equation to 2nd equation =.="

OpenStudy (mayankdevnani):

no,its not easy....

OpenStudy (anonymous):

its the other way round sir xD

OpenStudy (mayankdevnani):

sir ??? who is sir here ?

OpenStudy (mayankdevnani):

i am not a sir

OpenStudy (mayankdevnani):

just tell me mayank

OpenStudy (anonymous):

sorry.. okay

OpenStudy (mayankdevnani):

k

OpenStudy (mayankdevnani):

can you show your work(textbook solution)? from this , i can tell you

OpenStudy (anonymous):

ok hold on

OpenStudy (anonymous):

tbh the textbook also nvr say much

OpenStudy (mayankdevnani):

what?

OpenStudy (mayankdevnani):

just click image of your solution and then attach it

OpenStudy (anonymous):

there